OpenStudy (anonymous):

Find The Product For A Fan And A Medal! 1. 2x(3x - 5) 2. (x + 3)(x + 2) 3. (2x + 1)(x + 3) 4. (x + 1)(x^2 + x + 1) 5. (x + 2)(x - 2) 6. (x + 5)^2

OpenStudy (anonymous):

1. 6x^2-10x 2. x^2+2x+3x+6= x^2+5x+6 3.2x^2+6x+x+3 = 2x^2+7x+3 4.x^3+x^2+x+x^2+x+1 = x^3+2x^2+2x+1 5.x^2-4 6.(x+5)(x+5)

OpenStudy (e.mccormick):

6 is incomplete, and it is better explain how rather than to just do.

OpenStudy (anonymous):

Thanks for help

OpenStudy (anonymous):

Can you explain 6 to me?

OpenStudy (e.mccormick):

Know how FOIL works? They did an example, sort of, for #2. It has other names, but FOIL is common.

OpenStudy (e.mccormick):

FOIL stands for First then Outer then Inner then Last. It is just a way to keep your plce when multiplying things. Let me rework number 2. Then you should know how to do the same thing to number 6. So it started with: \((x + 3)(x + 2) \) \(\color{red}{\text{First}}~ (\color{red}{x} + 3)(\color{red}{x} + 2) \) so \(x\cdot x = \color{red}{x^2}\) is part of the answer. \(\color{blue}{\text{Outer}}~(\color{blue}{x} + 3)(x + \color{blue}{2}) \) so \(x\cdot 2 = \color{blue}{2x}\) is part of the answer. \(\color{green}{\text{Inner}}~(x + \color{green}{3})(\color{green}{x} + 2) \) so \(3\cdot x = \color{green}{3x}\) is part of the answer. \(\color{orange}{\text{Last}}~(x + \color{orange}{3})(x +\color{orange}{2}) \) so \(3\cdot 2 = \color{orange}{6}\) is part of the answer. \(\color{red}{x^2}+ \color{blue}{2x}+ \color{green}{3x}+ \color{orange}{6}= x^2+5x+6 \) Why you do it in this order is simple. Keep the outer and inner next to each other. This makes them easier to add or subtract because they are side by side.

OpenStudy (e.mccormick):

Had to fix something there... and the site is laggy so I don't know how long the small mistake will stay around.